The price of a condominium is ?$150,000. The bank requires a? 5% down payment and one point at the time of closing. The cost of the condominium is financed with a? 30-year fixed-rate mortgage at 7?%. down payment? mortgage? monthly payment?
Answer #1

Down payment = 5% of price of condominium = 5% x $150,000 = $7,500

Loan Amount (mortage) = Price - Down payment = $150,000 - $7,500 = $142,500

Monthly payments can be computed as follows -

Monthly Payments = Loan amount / PVIFA (rate, time periods)

PVIFA is present value interest factor annuity of $1 and can be computed using the following formula -

PVIFA = \frac{1-[\frac{1}{(1+r)^{n}}]}{r}

Since we have monthly payments, we require monthly rate and no. of months as time period.

Monthly rate (r) = 7% / 12 = 0.58333333%, n = No. of months = 30 x 12 = 360

PVIFA = \frac{1-[\frac{1}{(1+0.0058333333)^{360}}]}{0.0058333333} =150.307568556

Therefore, Monthly payment = $142,500 / PVIFA (0.58333333%, 360)

or, Monthly payment = $142,500 / 150.307568556 = $948.05605179 or $948.06